OpenStudy (anonymous):

FAN and MEDAL Can someone explain how exactly I create the formulas for binary compounds? Like the general rules?...I don't get it!

OpenStudy (anonymous):

Binary compounds are the ones made by 2 chemicals elements only. However, to write the formulas, it is important to recognize their oxidation number. Generally, as you know the group number of an element, the group number will be the same as it oxidation number. This oxidation number will determine the number of atoms of the other element binding with it and the same applies to the other element.

OpenStudy (anonymous):

But there is an exception: for those elements from groups 5 to 7, you have to do 8-the number of the group to determine their oxidation number. Let me show us a few examples: LiF, the formula is like this because the oxidation number or simply nox of Lithium is 1 and that of Fluorine is also 1. Cross exchanging the noxes, litium should have 1 atom and fluorine also 1. MgCl2. The formula is like this because the nox of Mg is 2 and that of Cl is 8-7=1. Cross exchanging the noxes, Mg should have 1 atom and Cl 2 atoms. Ca3N2, the formula is like this, because the nox of Ca is 2 and that of N is 8-5=3. Cross exchanging the noxes, Ca should have 3 atoms and N, 2 atoms. When a reaction is ionic,the metal comes first because it is a cation and the non-metal comes secondly. Hope you understood.

OpenStudy (anonymous):

This really helped!! okay so for something lets say i combined potassium and chlorine... i would take 7 -1 = 6 so I would write it as KCl6???? That seems wrong... :/ idk if I did that right.... @Hoslos

OpenStudy (anonymous):

No. Remember, if the element is from group 5 to 8, you have to do 8- the group number, being its nox. Chlorine is in group 7, so its nox is 8-7=1. Exchanging noxes, K needs 1 atom and so does Chlorine.

OpenStudy (anonymous):

so then its just KCl @Hoslos

OpenStudy (anonymous):

so basically to find the subscript number I just take 8 - the group number.....as long as it's in groups 5A - 8A ?

OpenStudy (anonymous):

Correct. I am glad you understood.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Okay! I think I've got it! thank you so much! Are you this good with math ??

OpenStudy (anonymous):

@Hoslos

OpenStudy (anonymous):

Well, it depends on the question. Let me see. Just spit it out.

OpenStudy (anonymous):

lol ok...its in the math section... I'll tag u in it its alg2

OpenStudy (anonymous):

okay I just tagged u....so it should alert ya...